OpenStudy (anonymous):

Points J (5,2) and K (-2,-5) are two vertices of a isosceles triangle. If L is the third vertex and has a y-coordinate of 4, what is the x-coordinate of L?

OpenStudy (campbell_st):

well there seem to be 2 options for this question 1. the sides LJ = KL so make the vertex L =(x, 4) and use the distance formula the 2nd choice would be JK is one of the equal sides... I think the 1st option is the easier

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so let L be the point (x, 4) so use the distance formula LJ \[d_{LJ} = \sqrt{(x - 5)^2 +(4 - 2)^2}\] and LK \[d_{LK} = \sqrt{(x + 2)^2 +(4 + 5)^2}\] equate the distances... since the triangle is isosceles and sqaure both sides and you get \[(x -5)^2 +(4 -2)^2 = (x +2)^2 + (4+5)^2\] so now you can distribute, simplify then solve for x hope it makes sense

OpenStudy (anonymous):

hey can we use the pythagoream theorum instead because we don't use the distance formula

OpenStudy (anonymous):

??

OpenStudy (campbell_st):

the distance formula is pythagoras. theorem 

OpenStudy (campbell_st):

now looking at pythagoras  you can find LJ and LK in terms of pythagoras then solve for x that's all I did above.